An article reported that, in a study of a particular wafer inspection process, 356 dies were examined by an inspection probe and 221 of these passed the probe. Assuming a stable process, calculate a 95% (two-sided) confidence interval for the proportion of all dies that pass the probe. (Round your answers to three decimal places.)

................................ , ......................................
Answer #1

solution:

given

n=356

x=221

p=x/n

=221/356

=0.62078

E=zalpha/2*sqr(p*(1-p)/n

=1.96*0.02571

=0.05040

C.i=0.62078+/-0.05040

=(0.57037,0.67118)